OpenStudy (anonymous):

What is x ? 3^x+2 = 1/9

OpenStudy (anonymous):

oops it is \[3^{x+2}=\frac{1}{9}\] and since \[3^{-2}=\frac{1}{9}\] you get \[x+2=-2\] so \[x=-4\]

OpenStudy (anonymous):

ohhhh ok ok I see